# MAAT: Mamba Adaptive Anomaly Transformer with association discrepancy for time series

### v1

@@ 5.1 Baseline Results
Analysis of Performance Across Datasets Refer to Table 4 The performance of our model is comprehensively evaluated against AnomalyTrans and DCdetector, two state-of-the-art methods, across five benchmark datasets: MSL, SMAP, SWaT, PSM, and SMD. The results in Table 4 demonstrate that our model consistently outperforms or closely matches the best-performing methods across most metrics, establishing its superiority in anomaly detection tasks. On the MSL dataset, our model achieves competitive results, with an F1-score of 95.05%, slightly lower than DCdetector 96.60%, representing a 1.6% decrease. However, our model surpasses AnomalyTrans in several metrics, including Accuracy Acc, where it achieves 98.92%, compared to AnomalyTrans 98.69%. While DCdetector excels in precision-related metrics such as Aff-P 51.84% and R_A_P 91.64%, our model demonstrates robust recall metrics, achieving Aff-R 96.49% and R_A_R 90.97%, which are critical for detecting subtle anomalies in noisy environments. On the SMAP dataset, our model shows slightly lower performance compared to AnomalyTrans and DCdetector. For instance, our F1-score 96.29% is marginally lower than DCdetector 97.02%, representing a 0.74% decrease. Similarly, our V ROC 92.06% and V PR 90.70% lag behind AnomalyTrans 95.52% and 93.77%, respectively. These minor trade-offs highlight areas for improvement, particularly in precision-related metrics. Nevertheless, our model maintains strong recall metrics, such as Aff-R 94.27%, underscoring its ability to detect anomalies effectively. On the SWaT dataset, our model achieves the highest Aff-P 56.26%, surpassing both AnomalyTrans 53.03% and DCdetector 52.40%. Additionally, our model demonstrates superior recall metrics, achieving Aff-R 97.79% and R_A_R 98.14%, which are marginally higher than DCdetector 97.67% and 98.43%, respectively. Our V _ROC 98.17% and V_PR 95.02% also outperform both competitors, highlighting its robustness in handling multivariate interactions and complex temporal patterns. Our model achieves the best performance across nearly all metrics on the PSM dataset. It surpasses DCdetector in F1-score 98.32% vs. 97.94%, a 0.39% increase and Aff-P 55.53% in respect to 54.71% in Dcdetector, a 1.5% increase. Our model excels in recall-related metrics, achieving Aff-R 85.06%, R_A_R 94.15%, and R A_P 95.09%, significantly outperforming both AnomalyTrans and DCdetector. These results underscore our model ability to handle high-dimensional sensor data with significant noise. On the SMD dataset, our model achieves the highest scores across all metrics, setting a new benchmark for performance. It surpasses AnomalyTrans in F1-score 92.30% vs. 87.18%, a 5.8% increase and Aff-P 59.19% vs. 54.36%, a 8.9% increase. Additionally, our model demonstrates superior recall metrics, achieving Aff-R 93.51%, R A_R 79.11%, and R_ A_P 75.41%, significantly outperforming both competitors. These gains highlight our model effectiveness in capturing intricate temporal patterns in highly non-stationary multivariate time series.
@@
In this section, we will analyze its performance specifically on the NIPS-TS-SWAN and NIPS-TS-GECCO datasets, as detailed in Table 5. Our method demonstrates consistent superiority across both NIPS-TS-SWAN and NIPS-TS-GECCO benchmarks. On NIPS-TS-SWAN, our model achieves 86.10% accuracy, outperforming AnomalyTrans by 1.53% and DCdetector by 0.16% . For precision, we attain 95.93%, surpassing AnomalyTrans by 5.22% and DCdetector by 0.45%. The most notable gains emerge in recall, where our method improves by 12.48% over AnomalyTrans and 0.36% over DCdetector. These improvements translate to a 73.76% F1-score, exceeding AnomalyTrans and DCdetector by 11.47% and 0.41%, respectively. On NIPS-TS-GECCO, our model achieves 98.68% accuracy, exceeding AnomalyTrans by 0.65% and DCdetector by 0.12%. Precision improves dramatically by 16.76% over AnomalyTrans and 4.16% over DCdetector, while recall rises by 41.52% and 10.27%, respectively. The F1-score reflects these gains, improving by 23.73% over AnomalyTrans and 6.19% over DCdetector.
Affinity metrics further highlight our method robustness: on NIPS-TS-GECCO, Aff-P improves by 7.88% over AnomalyTrans, while Aff-R increases by 11.89%. Validation ROC underscores this trend, with gains of 1.55% and 11.42% on SWAN and GECCO, respectively, over AnomalyTrans. Collectively, these results validate our method ability to balance precision, recall, and affinity-aware performance, establishing it as a state-of-the-art solution for time-series anomaly detection.
@@ 6.1 Ablation Study Analysis
In this section, we introduce the ablation study of our model to highlight the importance of each block in MAAT and its impact on the baseline Anomaly Transformer model. The results of the ablation study are shown in Table 6 and Table 7. Across five datasets—SMD, MSL, SMAP, SWaT, and PSM—the integration of Sparse Attention (SA), Mamba-SSM, and Gated Attention demonstrates consistent improvements over the baseline. These components work synergistically to address the challenges posed by noise, non-stationarity, and complex temporal dependencies inherent in multivariate time series data.
@@
On the SMD dataset, characterized by highly non-stationary multivariate time series with intricate patterns and significant noise, the baseline Anomaly Transformer achieves a robust recall and moderate precision, yielding an F1-score in the low 90% range. However, its dense self-attention mechanism renders it susceptible to noise-induced false positives. When Sparse Attention is introduced, the model preserves its baseline precision but suffers a recall drop of approximately 2.4%, resulting in an F1 decrease of about 1.15%—indicating that the block-wise formulation may oversimplify inter-variable dependencies and miss subtle anomalies. In contrast, integrating the Mamba-SSM component enhances long-range dependency capture, yet it incurs a reduction in precision by roughly 1% and a recall drop of about 2.2%, leading to an overall F1 decline of approximately 1.57%. Notably, the full MAAT model, which fuses both Sparse Attention and Mamba-SSM through a dynamic Gated Attention mechanism (Eqs. 1516), achieves a net F1 improvement of nearly 2% over the baseline—driven by a modest precision gain of around 0.6% and a substantial recall boost of roughly 3.5%. This adaptive gating effectively balances the contributions of local feature emphasis and long-term pattern refinement, robustly suppressing noise while preserving critical anomaly signals in the complex SMD environment.
@@
For the MSL dataset, which consists of sensor and actuator data with high variability and transient noise, the baseline Anomaly Transformer achieves a strong recall but moderate precision, resulting in an F1-score of approximately 93.93%. However, it struggles with noise-induced false positives. The introduction of Sparse Attention mitigates this issue, leading to a precision improvement of around 0.7%, but at the cost of a 2% recall reduction, ultimately lowering the F1-score by about 1.1%. This trade-off arises as SA effectively filters noise but inadvertently excludes short-term critical patterns. Conversely, integrating Mamba-SSM enhances precision further by approximately 1.15% over the baseline while maintaining recall stability, yielding a slight F1 improvement of 0.12%. This enhancement is attributed to Mamba’s ability to stabilize reconstructions via linear long-range dependency modeling. The full MAAT model synthesizes these improvements through Gated Attention, achieving an F1-score gain of 1.12% over the baseline—driven by a recall increase of 2.3% while maintaining competitive precision. During periods of transient noise, the adaptive gating mechanism (Eqs. 15) prioritizes the denoised outputs from Mamba while preserving Sparse Attention’s localized anomaly signals. This balanced integration reinforces MAAT’s robustness to short-lived disturbances, effectively enhancing anomaly detection in highly dynamic environments.
@@
The SMAP dataset, which consists of satellite telemetry data with long-range dependencies and subtle anomalies, highlights MAAT’s capability to balance global context with localized precision. The baseline Anomaly Transformer achieves a near-perfect recall but exhibits overfitting, limiting precision to 93.59% and resulting in an F1-score of 96.41%. To address this, Sparse Attention reduces overfitting by sparsifying redundant global interactions (Eq. 8), but this comes at the cost of slightly weaker long-range anomaly detection. Integrating Mamba-SSM substantially improves precision by approximately 2%, leading to an overall F1-score improvement of 0.67%, as it selectively propagates long-term dependencies while filtering high-frequency noise. The full MAAT model harmonizes these strengths through Gated Attention (Eqs. 1516), achieving a final precision of 94.75%, recall of 99.33%, and F1-score of 96.99%, surpassing the baseline by 0.58%. This enhancement is driven by the gating mechanism, which dynamically prioritizes Mamba’s global context for slow-evolving anomalies while leveraging Sparse Attention’s local focus for abrupt deviations. This adaptability ensures MAAT effectively captures both transient and long-term anomalies, making it well-suited for SMAP’s unique challenges.
@@
The SWaT dataset, consisting of 51-dimensional sensor data from a water treatment system with intricate multivariate interactions, demonstrates the necessity of capturing inter-sensor dependencies for precise anomaly detection. The baseline Anomaly Transformer achieves a high recall of 99.41% but struggles with false positives, leading to a precision of 93.59% and an F1-score of 96.41%. Introducing Sparse Attention significantly degrades precision by 4.06%, reducing the F1-score to 93.96%, as it fails to model critical cross-sensor relationships. Conversely, integrating Mamba-SSM enhances recall to a perfect 100% and increases the F1-score to 96.55% by effectively capturing complex interdependencies while filtering redundant sensor noise. The full MAAT model strategically combines these mechanisms through Gated Attention (Eqs. 1516), achieving a final precision of 93.33%, recall of 100%, and F1-score of 96.50%. This result matches Mamba’s recall while mitigating Sparse Attention’s limitations, ensuring MAAT effectively balances precision and recall in SWaT’s multivariate anomaly detection setting. The PSM dataset, comprising 25-dimensional sensor data with significant noise, highlights the necessity of mitigating false positives while maintaining robust anomaly detection. The baseline Anomaly Transformer performs well, achieving an F1-score of 97.47%, but is affected by noise-induced misclassifications, leading to a recall of 97.81% and a precision of 97.14%. Sparse Attention enhances recall by 1.56%, reaching 99.37%, which raises the F1-score to 98.31%, though it slightly reduces precision by 0.13% due to its localized attention windows (Eq. 8) focusing on critical time steps. Conversely, Mamba-SSM improves precision to 97.41% while stabilizing reconstructions, resulting in an F1-score of 97.89%. The full MAAT model integrates these strengths via Gated Attention (Eqs. 1516), achieving a balanced precision of 97.48%, recall of 99.17%, and F1-score of 98.32%. This adaptability is crucial in PSM’s noisy, high-dimensional setting, where rigid attention mechanisms struggle to isolate genuine anomalies. Furthermore, the Association Discrepancy metric (Eq. 17) refines anomaly detection by penalizing deviations from expected sensor dependencies, reducing false positives. These findings reinforce MAAT’s effectiveness in dynamically balancing precision and recall, ensuring robustness across diverse time-series anomaly detection tasks.
@@
The NIPS-TS-GECCO dataset, a benchmark for drinking water quality with sporadic, high-frequency anomalies and IoT sensor noise, highlights MAAT’s strengths and limitations. The baseline Anomaly Transformer (AnomalyTrans) struggles, achieving low precision ($P=25.65\%$) and recall ($R=28.48\%$, $F1=29.09\%$), as its dense self-attention conflates short-lived anomalies with pervasive noise.
Introducing Sparse Attention (AnomTr+SA) improves performance ($P=45.19\%$, $R=84.25\%$, $F1=58.82\%$) by localizing attention to critical time steps via $\Omega_{i}$, which defines a local window (Eq. 12). However, its rigid block size ($block\_size=100$) limits recall for variable-duration anomalies.
AnomTr+Mamba further refines results ($P=36.00\%$, $R=51.78\%$, $F1=42.47\%$) using Mamba-SSM’s state-space model (Eq. 10) for denoising. Yet, its linear dependency modeling underperforms for abrupt, short-term anomalies.
The full MAAT model integrates these components via Gated Attention, achieving $P=42.41\%$, $R=70.00\%$, $F1=52.82\%$, a 13.3% F1 gain over DCdetector ($F1=46.6\%$). The gating mechanism (Eq. 15–16) adaptively prioritizes Sparse Attention during sudden events ($g\to 0$) and Mamba’s context during stable periods ($g\to 1$). However, the Association Discrepancy (Eq. 3) is less effective here due to minimal dependency shifts in momentary spikes.
While MAAT outperforms alternatives, its precision trails IForest ($43.9\%$ vs. $42.41\%$). This trade-off reflects MAAT’s focus on higher recall ($70.00\%$ vs. IForest’s $35.3\%$), making it ideal for real-world scenarios where missing critical anomalies is riskier than false alarms.
@@
The NIPS-TS-SWAN dataset, derived from solar photospheric vector magnetograms, focuses on detecting subtle anomalies in high-dimensional, temporally complex solar activity patterns. The baseline Anomaly Transformer (AnomalyTrans) achieves moderate performance ($P=90.71\%$, $R=47.43\%$, $F1=62.29\%$), as dense self-attention struggles to isolate rare solar anomalies (e.g., magnetic field instabilities) from noisy, non-stationary signals.
AnomTr+SA improves precision ($P=97.03\%$, $F1=73.60\%$) by leveraging Sparse Attention (Eq. 12) to focus on localized time windows critical for transient events (e.g., solar flares). However, its rigid block size ($block\_size=100$) limits recall ($R=59.28\%$) by truncating long-range magnetic field dependencies.
AnomTr+Mamba enhances recall ($R=59.34\%$, $F1=73.65\%$) using Mamba-SSM’s state-space model (Eq. 10), which captures gradual solar cycles and magnetic shifts. While Mamba improves global context, its linear dynamics underperform for abrupt anomalies (e.g., sudden polarity reversals), slightly reducing precision ($P=97.06\%$).
The full MAAT model integrates these strengths via Gated Attention, achieving $P=95.93\%$, $R=59.91\%$, $F1=73.76\%$—a 18.5% F1 gain over AnomalyTrans and 0.55% over DCdetector. The gating mechanism (Eq. 15–16) adaptively prioritizes Mamba’s long-term dynamics ($g\to 1$) during stable periods (e.g., gradual magnetic buildup) and SA’s localized focus ($g\to 0$) during rapid anomalies (e.g., coronal mass ejections).
Association Discrepancy (Eq. 3) enhances detection by penalizing deviations between prior-association (expected solar magnetic correlations) and series-association (observed sparse dependencies). For instance, anomalies in magnetic flux tube alignments are flagged when sparse attention detects localized disruptions inconsistent with Gaussian-kernel expectations.
@@
Despite its gains, MAAT’s recall ($59.91\%$) slightly trails IForest ($59.8\%$), reflecting a precision-recall trade-off: MAAT prioritizes high-confidence anomalies (e.g., pre-eruption magnetic shearing) over marginal signals, reducing false alarms in noisy solar data. This aligns with real-world solar monitoring, where minimizing false positives is critical to avoid unnecessary alerts.
@@
Table 7: Ablation study of precision, recall, and F-score on NIPS datasets. “SA” stands for Sparse Attention; the best metrics are in bold.

### v2

@@ 5.1 Baseline Results
Analysis of Performance Across Datasets Refer to Table 3 The performance of our model is comprehensively evaluated against AnomalyTrans and DCdetector, two state-of-the-art methods, across five benchmark datasets: MSL, SMAP, SWaT, PSM, and SMD. The results in Table 3 demonstrate that our model consistently outperforms or closely matches the best-performing methods across most metrics, establishing its superiority in anomaly detection tasks.
On the MSL dataset, our model surpasses AnomalyTrans in several metrics, including Accuracy Acc, where it achieves 98.92%, compared to AnomalyTrans’s 98.69%. While DCdetector excels in precision-related metrics such as Aff-P 51.84% and R_A_P 91.64%, our model demonstrates robust recall metrics, achieving Aff-R 96.49% and R_A_R 90.97%, which are critical for detecting subtle anomalies in noisy environments.
On the SMAP dataset, our model shows slightly lower performance compared to AnomalyTrans and DCdetector. For instance, our V ROC 92.06% and V PR 90.70% lag behind AnomalyTrans 95.52% and 93.77%, respectively. These minor trade-offs highlight areas for improvement, particularly in precision-related metrics. Nevertheless, our model maintains strong recall metrics, such as Aff-R 94.27%, underscoring its ability to detect anomalies effectively.
On the SWaT dataset, our model achieves the highest Aff-P 56.26%, surpassing both AnomalyTrans 53.03% and DCdetector 52.40%. Additionally, our model demonstrates superior recall metrics, achieving Aff-R 97.79% and R_A_R 98.14%, which are marginally higher than DCdetector 97.67% and 98.43%, respectively. Our V _ROC 98.17% and V_PR 95.02% also outperform both competitors, highlighting its robustness in handling multivariate interactions and complex temporal patterns.
Our model achieves the best performance across nearly all metrics on the PSM dataset, a 1.5% increase in Aff-P vs. DCdetector, our model excels in recall-related metrics, achieving Aff-R 85.06%, R_A_R 94.15%, and R A_P 95.09%, outperforming AnomalyTrans and DCdetector. These results underscore our model’s ability to handle high-dimensional sensor data with significant noise.
On the SMD dataset, our model achieves the highest scores across all metrics, setting a new benchmark for performance. It surpasses AnomalyTrans in 5.8% and 8.9% increase in F1-score and Aff-P respectively. Additionally, our model demonstrates superior recall metrics, achieving Aff-R 93.51%, R A_R 79.11%, and R_ A_P 75.41%, significantly outperforming both competitors. These gains highlight our model’s effectiveness in capturing intricate temporal patterns in highly non-stationary multivariate time series.
@@
In this section, we will analyze its performance specifically on the NIPS-TS-SWAN and NIPS-TS-GECCO datasets, as detailed in Table 4. Our method demonstrates consistent superiority across both NIPS-TS-SWAN and NIPS-TS-GECCO benchmarks outperforming accuracies of AnomalyTrans by 1.53% and DCdetector by 0.16%. For precision, we surpassed AnomalyTransformer by 5.22% and DCdetector by 0.45%. The most notable gains emerge in recall; method improves by 12.48% over AnomalyTransformer and 0.36% over DCdetector. MAAT’s F1-score, exceeded AnomalyTransformer and DCdetector by 11.47% and 0.41%, respectively. On NIPS-TS-GECCO, our model’s Accuracy surpassed AnomalyTransformer by 0.65% and DCdetector by 0.12%. For precision, we had an increase of 16.76% compared to AnomalyTransformer and 4.16% over DCdetector. The recall shows significant gains, improving by 41.52% and 10.27%, respectively. Furthermore, the F1-score demonstrates these advancements, rising by 23.73% over AnomalyTransformer and 6.19% over DCdetector.
Affinity metrics further emphasize the robustness of our method: on the NIPS-TS-GECCO dataset, Aff-P shows an improvement of 7.88% over AnomalyTrans, while Aff-R increases by 11.89%. The validation ROC corroborates this trend, revealing gains of 1.55% and 11.42% on SWAN and GECCO, respectively, compared to AnomalyTrans. Together, these results validate our method’s capability to effectively balance precision, recall, and Recallllinity-aware performance, positioning it as a state-of-the-art solution for time-series anomaly detection.
@@ 6.1 Ablation Study Analysis
This section, presents the results of the ablation study of our model, emphasizing the significance of each component within MAAT and their effects on the baseline Anomaly Transformer model. The findings from the ablation study are summarized in Table 5 and Table 6. Across five datasets SMD, MSL, SMAP, SWaT, and PSM the incorporation of Sparse Attention (SA), Mamba-SSM, and Gated Attention consistently shows enhancements over the baseline.
@@
The baseline Anomaly Transformer demonstrates strong recall and moderate precision on SMD dataset. However, its dense self-attention mechanism makes it vulnerable to false positives due to noise. The integration of Sparse Attention allows it to maintain baseline precision but leads to a recall decrease of approximately 2.4%, resulting in an F1-score reduction of about 1.15%. This suggests that the block-wise formulation may oversimplify inter-variable dependencies and overlook subtle anomalies. Conversely, the integration of the Mamba-SSM component enhances the model’s ability to capture long-range dependencies, albeit at the cost of a precision reduction of approximately 1% and a recall decrease of about 2.2%, culminating in an overall F1 decline of approximately 1.57%. Remarkably, the complete MAAT model, which combines Sparse Attention and Mamba-SSM through a dynamic Gated Attention mechanism (Eqs. 78), achieves a net F1 improvement of nearly 2% over the baseline, fueled by a modest precision gain of around 0.6% and a significant recall enhancement of roughly 3.5%. This adaptive gating effectively balances the contributions of precision and long-term pattern refinement, robustly mitigating noise while preserving critical anomaly signals within the complex SMD environment.
@@
For the MSL dataset, which comprises sensor and actuator data characterized by high variability and transient noise, the baseline Anomaly Transformer demonstrates strong recall but moderate precision, resulting in an F1-score of approximately 93.93%. However, it encounters challenges with false positives induced by noise. The introduction of Sparse Attention alleviates this issue, yielding a precision improvement of around 0.7%, albeit at the expense of a 2% reduction in recall, which ultimately lowers the F1-score by about 1.1%. This trade-off occurs because Sparse Attention effectively filters out noise but inadvertently excludes crucial short-term patterns. In contrast, the integration of Mamba-SSM enhances precision by approximately 1.15% over the baseline while stabilizing recall, resulting in a modest F1 improvement of 0.12%. This enhancement is attributed to the stabilization of reconstructions through linear long-range dependency modeling. The comprehensive MAAT model synthesizes these advancements through Gated Attention, achieving a gain of 1.12% in F1-score over the baseline, driven by a recall increase of 2.3% while maintaining competitive precision. During instances of transient noise, the adaptive gating mechanism (Eqs. 7) emphasizes the denoised outputs of Mamba while preserving the localized anomaly signals from Sparse Attention. This balanced integration strengthens MAAT’s robustness against short-lived disturbances, significantly enhancing anomaly detection in highly dynamic environments.
@@
The SMAP dataset, which comprises satellite telemetry data characterized by long-range dependencies and subtle anomalies, underscores MAAT’s ability to balance global context with localized anomaly detection. The Anomaly Transformer achieves an almost perfect recall but suffers from overfitting, which limits its precision to 93.59%, resulting in an F1-score of 96.41%. To mitigate this issue, Sparse Attention addresses overfitting by reducing redundant global interactions (see Eq. 2), although this leads to slightly weaker long-range anomaly detection. The integration of Mamba-SSM significantly enhances precision by approximately 2%, yielding an overall F1-score improvement of 0.67%. Mamba-SSM accomplishes this by selectively propagating long-term dependencies while filtering out high-frequency noise. The complete MAAT model harmonizes these strengths through Gated Attention (refer to Eqs. 78), achieving a final precision of 94.75%, a recall of Y%, and an F1-score of 96.99%, exceeding the baseline by 0.58%. This enhancement is propelled by the gating mechanism, which prioritizes Mamba’s global context for slow-evolving anomalies while leveraging Sparse Attention’s local focus for abrupt changes. This adaptability enables MAAT to effectively capture both transient and long-term anomalies, making it particularly well-suited to address the unique challenges presented by the SMAP dataset.
@@
The SWaT dataset, which includes 51-dimensional sensor data from a complex water treatment system, highlights the importance of capturing inter-sensor dependencies for effective anomaly detection. The baseline Anomaly Transformer achieves an impressive recall of 99.41%; however, it faces challenges with false positives, resulting in a precision of 93.59% and an F1-score of 96.41%. The introduction of Sparse Attention significantly reduces precision by 4.06%, causing the F1-score to drop to 93.96%, as this approach fails to model critical cross-sensor relationships adequately. In contrast, the integration of Mamba-SSM improves recall to an exceptional 100% and raises the F1-score to 96.55% by effectively capturing complex interdependencies while filtering out redundant sensor noise. The complete MAAT model strategically combines these mechanisms through Gated Attention (Eqs. 78), achieving a final precision of 93.33%, along with its aforementioned recall and an F1-score of 96.50%. This development underscores Mamba’s ability to recognize the limitations of sparse attention, ensuring that MAAT maintains a balanced approach to precision and recall in a multivariate anomaly detection context. The PSM dataset, which consists of 25-dimensional sensor data characterized by substantial noise, underscores the importance of reducing false positives while ensuring effective anomaly detection. The baseline Anomaly Transformer exhibits commendable performance, attaining an F1-score of 97.47%. However, it suffers from noise-induced misclassifications, reflected in a recall of 97.81% and precision. In contrast, Sparse Attention improves recall to 99.37%, consequently elevating the F1-score to 98.31%. This enhancement comes at the cost of a slight reduction in precision, down by 0.13%, due to its localized attention windows (Eq. 2), which focus on critical time steps. On the other hand, Mamba-SSM enhances precision to 97.41% while stabilizing reconstructions, leading to an F1-score of 97.89%. The comprehensive MAAT model combines these strengths through Gated Attention (Eqs. 78), resulting in a balanced precision of 97.48%, an improved recall, and an F1-score of 98.32%. This adaptability is essential in the noisy, high-dimensional context of PSM, where traditional attention mechanisms often struggle to identify true anomalies. Additionally, the Association Discrepancy metric (Eq. 9) refines anomaly detection by penalizing deviations from expected sensor dependencies, thereby reducing false positives. These findings reinforce the effectiveness of MAAT in dynamically balancing precision, recall, and robustness across a range of time-series anomaly detection tasks.
@@
The NIPS-TS-GECCO dataset is a benchmark for drinking water quality, characterized by sporadic, high-frequency anomalies and noise from IoT sensors. This dataset highlights the limitations of the baseline Anomaly Transformer, which attains an F1-score of only 29.09%. In contrast, implementing Sparse Attention (AnomTr+SA) yields significant improvements, with approximately 19.5% in Precision, 55.8% in Recall, and an increase in the F1-score. These enhancements are achieved by focusing on critical time steps using $\Omega_{i}$ (Eq.2), effectively reducing the conflation of transient anomalies with sensor noise. Conversely, incorporating Mamba-SSM (AnomTr+mamba) results in more modest gains, with an increase of around 10.4% in Precision, 23.3% in Recall, and a corresponding improvement in the F1-score, thanks to its state-space model for denoising. The linear dependency modeling is less effective in addressing abrupt, short-term anomalies. However, when both components are integrated using Gated Attention (AnomTr+mamba+SA), the comprehensive MAAT model demonstrates significant improvements, achieving approximately 16.8% in Precision, 41.5% in Recall, and an increase in the F1-score compared to the baseline. This results in an F1-score that is 13.3% higher than the DCdetector. The gating mechanism prioritizes Sparse Attention during sudden events (where $g\to 0$) while leveraging Mamba’s contextual modeling during stable periods (where $g\to 1$). In this context, Association Discrepancy’s effectiveness is reduced due to minimal dependency shifts during brief spikes.
@@
The NIPS-TS-SWAN dataset, which is derived from solar photospheric vector magnetograms, concentrates on identifying subtle anomalies within high-dimensional, temporally complex solar activity patterns. The baseline Anomaly Transformer (AnomalyTrans) demonstrates moderate performance, achieving a precision of 90.71%, a recall value of Recall%, and an F1-score of 62.29%. Its dense self-attention mechanism, however, struggles to distinguish rare solar anomalies from noisy, non-stationary signals. By employing Sparse Attention, which hones in on localized time windows crucial for transient events, the AnomTr+SA variant improves precision by approximately 6.3% percentage points, recall by Recall percentage points, and F1-score by 11.3% percentage points compared to the baseline.
Conversely, the integration of Mamba-SSM into AnomTr+mamba enhances global context, resulting in improvements of about +6.4%, 11.9%, and 11.4% in precision, recall, and F1-score, respectively, although its linear dynamics tend to underperform for sudden anomalies such as abrupt polarity reversals. Remarkably, when both components are combined using Gated Attention, the complete MAAT model achieves net enhancements of approximately 5.2% in precision, 12.5% in recall, and Recall percentage points in F1-score relative to the baseline—amounting to an 18.5% relative gain in F1 over AnomalyTrans and a 0.55% improvement over DCdetector. The gating mechanism adaptively prioritizes Mamba’s long-term dynamics during stable periods ($g\to 1$) and Sparse Attention’s localized focus during rapid anomalies ($g\to 0$). Moreover, the Association Discrepancy metric (Eq. 3) refines anomaly detection by penalizing deviations between expected solar magnetic correlations and observed sparse dependencies, effectively identifying anomalies in magnetic flux tube alignments.
@@
Despite these advancements, the recall of MAAT remains marginally lower than that of IForest, highlighting a precision-recall trade-off where the model favors high-confidence anomalies over marginal signals—a vital consideration in real-world solar monitoring, where minimizing false positives is critical.
The NIPS-TS-SWAN dataset, derived from solar photospheric vector magnetograms, focuses on detecting subtle anomalies within high-dimensional and temporally complex patterns of solar activity. The baseline Anomaly Transformer (AnomalyTrans) demonstrates moderate performance, achieving a precision of 90.71%, and an F1-score of 62.29%. However, its dense self-attention mechanism has difficulty distinguishing rare solar anomalies from noisy, non-stationary signals. By utilizing Sparse Attention to focus on localized time windows critical for transient events, the AnomTr+SA variant achieves improvements of approximately +6.3% in precision, +11.9% in recall, and an increase in the F1-score compared to the baseline. Conversely, integrating Mamba-SSM in AnomTr+mamba enhances global context, resulting in gains of roughly +6.4% in precision and +11.9% in recall, along with an increase in the F1-score.
However, the linear dynamics tend to perform less effectively during abrupt anomalies, such as sudden polarity reversals. Importantly, when these two components are integrated using Gated Attention, the full MAAT model exhibits net gains of approximately +5.2% in Precision, +12.5% in Recall, and notable improvements in F1-score compared to the baseline. This results in an 18.5% relative gain in F1 over AnomalyTrans and a 0.55% increase in performance over the DCdetector. The gating mechanism allows for adaptive focus, prioritizing Mamba’s long-term dynamics during stable periods (when $g\to 1$) and emphasizing Sparse Attention’s localized responses during rapid anomalies (when $g\to 0$).
The Association Discrepancy metric enhances anomaly detection by penalizing the differences between expected solar magnetic correlations and the observed sparse dependencies. This effectively identifies inconsistencies in the alignments of magnetic flux tubes. However, despite these improvements, MAAT’s recall rate is still slightly lower than IForest’s. This demonstrates a precision-recall trade-off, in which the model focuses on high-confidence anomalies rather than weaker signals. This is a critical consideration for real-world solar monitoring, where it is essential to minimize false positives.
@@
Table 6: Ablation study Precisionion, recall,Recallllcore on NIPS datasets. “SA” stands for Sparse Attention; the best metrics are in bold.